# Light Transport-aware Diffusion Posterior Sampling for Single-View Reconstruction of 3D Volumes

### v2

@@ Abstract
Figure 1: Given a single view ($y$) of a volume ($GT$), we reconstruct a volume ($\hat{V}$) that matches $y$ under the same lighting conditions, resulting in a synthesized view ($\hat{y}$). Differentiable volume rendering optimizes physical volume parameters using Stochastic Gradient Descent (SGD) in parameter space while simultaneously performing posterior sampling, conditioned on the observation, in the latent space of a trained diffusion model (blue curve). Ambiguities due to the absence of information about unseen parts of the volume are reduced by gradually steering the reverse diffusion process toward the most plausible reconstruction under the given view.
@@ 1 Introduction
Differentiable rendering (DR) uses a rendering process to recover the shape and optical material parameters from images [25, 70, 37, 11]. DR enables the calculation of gradients of rendered images with respect to these parameters, including position, texture, lighting, shape, and other attributes. The challenge increases significantly when these parameters describe complex distributions of volumetric materials, such as clouds, smoke, or fire. In such scenarios, the problem becomes so ill-posed that it requires dozens, if not hundreds, of different views to adequately constrain the object parameters  [44, 45, 27]. It is now widely accepted that reconstructing the internal density distribution of highly dense volumes is nearly impossible due to the high uncertainty in the light scattering process and the presence of vanishing gradients. This limitation can only be alleviated by incorporating prior information during reconstruction.
@@
Our proposed approach addresses these challenges by employing a diffusion prior to guide a Physically-based Differentiable Volume Renderer (PDVR) toward reconstructing a plausible volumetric field. The prior is trained to learn the probability distribution of data conditioned on noisy versions of that data. In contrast to previous approaches, our approach includes controlled variations in the diffusion step by considering the image gradient with respect to the learned latent space representation. This approach steers the reconstruction toward a realistic 3D density distribution, ensuring that the generated structure aligns well with observed data and maintains realistic spatial consistency
@@
  • A novel parametric Diffusion Posterior Sampling (DPS) technique utilizing a shape-centric prior with a physically-based differentiable volume renderer.
@@ 2 Related Work
NeRFs have been generated initially with MLP-based Scene Representation Networks (SRNs) [55], which have later been used to compactly encode volumetric scalar fields using the emission-absorption optical model  [33, 74]. Alternative to the use of SRNs, adversarial approaches have recently emerged. They use 2D images to stochastically condition the 3D reconstruction using an adversarial loss [53, 4, 16, 43, 87]. In this context, sparse tri-plane volumetric models have been proposed to reduce the memory consumption at improved training efficiency of NeRFs [14, 5]. While NeRF-based approaches usually assume that images of the scene from many different viewpoints exist, recent advancements have shown their potential to also perform single-view reconstruction [4, 39, 53, 69, 53, 39, 76, 29].
@@
Zhang et al.  [80] present a general framework, including volumetric media, for calculating radiance derivatives with respect to changes of scene parameters. This framework has later been extended to make it applicable to path tracing including random sampling [81]. Properties of the differentiation of integrators are analyzed by Zeltner and Monte [77]. Forward mode automatic differentiation [44] for differentiable rendering is nowadays replaced by radiative backpropagation [45] to decrease the required memory, yet at the expense of multiple branches along light paths and quadratic time complexity thereof. Performance increases are achieved by reusing radiances along light paths [67], and by avoiding recursive radiance estimates at scattering locations with dedicated sampling methods for estimating derivatives of volumetric scattering  [46]. For multi-view reconstruction of volumetric fields in the presence of global light transport, singular path sampling in combination with in-scattering relaxation and an exponential moving average shows improved reconstruction fidelity [27]. Under the assumption of an emission-absorption optical model, the ”inversion trick” enables fast automatic differentiation for volume reconstruction and transfer function learning  [73]. Physical constraints are combined with self-supervision for the reconstruction of single-scattering flow fields from single-view videos [12].
@@ 4 Method
Figure 2: Cloudy Dataset: Photorealistic renderings of randomly selected clouds from our dataset, illustrating natural variations and details.
@@ 4.1 Cloudy - a 3D Clouds Dataset
To add natural randomness and represent diverse distributions of warm columns to the clouds, we apply Perlin noise functions and varied particle emission shapes. Figure 2 shows a random selection of clouds from our dataset, which are rendered under different lighting conditions. The density fields are numerically simulated on regular 3D grids at a resolution of approximately $512\times 256\times 512$.
@@ 4.2 Volume Latent Encoding
We introduce an implicit neural representation for a volume $\mathcal{V}$ based on a single projection, which we refer to as monoplanar. Unlike previous approaches that use positional feature embeddings like triplane or tensor decomposition, our method involves sampling a window across a single projected axis, centered at the coordinate of interest, to extract the final features.
Let $G:\mathbb{R}^{2}\rightarrow\mathbb{R}^{N}$ be a continuous two-dimensional field of features. Define $F(x,y,z)$ as a piecewise linear function with respect to $z$, specified over a regular grid in the interval $[-1,1]$ by the values $G_{i}(x,y)$. For a given position $(x,y,z)$, the embedded feature vector $f_{k}$ is computed as
$f_{k}(x,y,z)=F(x,y,z+\Delta_{k}),$
where $\Delta_{k}$ is a linearly spaced offset within the interval $[-1,1]$. Shifted values that fall outside this interval are clamped. Finally, a filter implemented with an MLP is applied to produce the final value:
$\mathcal{V}(\cdot)=\textsc{MLP}(f_{1}(\cdot),...,f_{k}(\cdot)).$
@@
In practice, we parameterize $G$ with a coarse grid of size $128\times 128$ and 32 features. A convolutional upsampler is applied to increase the resolution to $256\times 256\times 64$. Once upsampled, the feature vectors at a specific grid position $(x_{0},y_{0})$ are obtained using bicubic interpolation (Figure 3).
Figure 4: Diffusion based cloud synthesis. Clouds generated with our diffusion model, demonstrating convincing appearance under realistic lighting conditions and physical parameters.
@@ 4.3 Volume Latent Space
Since cloud formations are invariant to arbitrary rotations and minor scaling along the $xy$-plane, we apply 14 such operations to the clouds and augment the dataset by these instances. The analog transformations are applied to the latent codes as an initial solution, which is then subsequently refined via optimization. While the transformed latent already represents a plausible volume, the refinement prevents the diffuser from learning patterns that emerge purely from resampling, i.e., due to boundaries and clamping (see the supplementary material for an example). Including the $8$ invariant transformations (flips and transposes), we obtain a total of $1,000\times 14\times 8$ volume instances for training. Figure 4 demonstrates the effectiveness of our diffuser in generating new, unconditional volumetric instances. The ability to produce clouds with realistic shape and interior is demonstrated in Figure 5.
Figure 5: Diffusion Sampling. First column: A cloud from the Cloudy dataset. Subsequent columns show clouds generated by our diffusion model. First two rows show the clouds under different lighting conditions, demonstrating realistic cloud-light interaction. Bottom row shows cross-sectional slices through the volumes, demonstrating realistic interiors of diffused clouds.
@@ 4.4 Parameterized Posterior Sampling
Let us now assume that a proper posterior sampling method $p(\theta|y;\phi)$ is available, meaning that given an observation $y$ and a forward model $y=\mathcal{A}(\theta;\phi)+\eta$, we can draw samples $\theta$ that satisfy the observation. In our case, $\mathcal{A}(\theta;\phi)$ encapsulates both the decoding of the volume from $\theta$ and the rendering depending on $\phi$, i.e.,
$\mathcal{A}(\theta;\phi):=\mathcal{R}(\mathcal{D}(\theta),\phi).$
@@
After determining $\hat{\phi}$, the final latent representation $\theta$ can be sampled from the posterior distribution $\theta\sim p(\theta|y;\hat{\phi})$.
@@ 4.5 Optimization
Depending on the complexity of $\mathcal{A}(\theta;\phi)$ with respect to the parameters, it may be advantageous to reuse the same sample $\theta$ for multiple steps during the overall optimization. This strategy reduces the need for repeated sampling, while still allowing effective updates to $\phi$ over several iterations. This can be particularly useful when $\mathcal{A}(\theta;\phi)$ involves expensive operations or when the gradient propagation is computationally intensive.
@@
Finally, an optional refinement step can be applied, which enforces data consistency [57] before the latent $\theta$ is reused to improve $\phi$ and diffuse for the next step. This is achieved by directly optimizing the latent without any prior supervision. The rationale is that certain features will be preserved, allowing the latent to converge more quickly without constraints. Additionally, if ambiguity arises, it is advantageous for it to be reflected initially in the parameter that is subsequently “cleaned” by the prior.
The proposed optimization is outlined in Algorithm 1.
@@
$\hat{\phi}=\operatorname*{arg\,min}_{\phi}\mathbb{E}_{p(\theta|y;\phi)}\left[\|y-\mathcal{R}(\mathcal{D}(\theta),\phi)\|^{2}_{2}\right]$
$\theta\sim p(\theta\,|\,y;\hat{\phi})$
@@
return $\phi_{N},\theta_{N}$
@@ 5 Results
In this section, we demonstrate the effectiveness of our method for different use cases. All modules are implemented in Pytorch [48] and Vulkan SDK. Further details are provided in the supplementary material. The code and the Cloudy dataset will be released upon acceptance.
@@ 5.1 Diffusion Posterior Sampling
In a first experiment we shed light on the potential of DPS for single-view volume reconstruction. Through this experiment, we do not optimize for any physical parameters affecting the cloud appearance, but solely assess the strength of the volume diffusion prior when used to constrain the differentiable volume renderer.
@@ 5.2 Monoplanar Representation
To access the quality that is achieved with the proposed monoplanar latent representation, we perform a series of experiments with the monoplanar, triplanar and dense grid representations. Table  1 shows the average values for each metric across nine reconstructions using clouds from the Cloudy dataset.
@@
While PSNR, RMSE and MAE consider the full volume at $256\times 128\times 256$ resolution, SSIM [71] considers the center slice. Our proposed monoplanar representation quantitatively outperforms the other state-of-the-art representations in terms of reconstruction fidelity.
@@
5.3 Cloud Recovery from Transmittance Measures
5.4 Comparative Evaluation
To compare our novel DPS approach with previous methods for reconstructing 3D clouds from images, we evaluate DPS alongside Differentiable Ratio-Tracking (DRT) [46] and Singular Path Sampling (SPS) [27]. Since both DRT and SPS require multiple views to achieve accurate results, we use three images taken from below a cloud for the reconstructions.
We evaluate DPS under three different settings: (1) using only a single view (DPS1), (2) using all three views (DPS3), and (3) performing three restarts of the diffusion from a noisy version of a previously reconstructed latent (DPS3x3). The last setting aligns with diffuse-denoise strategies, progressively adjusting the initial noise toward the observed data to improve guidance stability. Results are shown in Fig. 9.
Figure 9: Reconstruction comparison using two distinct cloud configurations. In each experiment, the three leftmost $2\times 2$ images depict views from below the cloud and are used for reconstruction. The fourth image serves as a test view from above the cloud. DPS1 utilizes only one image for reconstruction, whereas all other methods use all three images. The reconstruction time is reported in minutes (top), along with the LPIPS [82] metric value (bottom), which quantifies the perceptual similarity between the ground truth and the synthesized views.
5.5 Recovering Light Conditions
Figure 10: Recovering $\phi$. Top: Reconstructions using parameterized DPS under three scenarios—when no physical parameters are unknown (None), when the background radiance is unknown (Background), and when the entire lighting condition is unknown (Environment). Bottom: Evolution of the recovered background (top) and environment (bottom). Final column shows the lighting condition used to render the test views.
@@ 5.6 Super-Resolution
Figure 11: Cloud Super-Resolution. From a cloud on a $32\times 16\times 32$ grid, the diffuser reconstructs a density distribution on a $256\times 128\times 256$ grid. This process adds fine details and internal structures, demonstrating the model’s ability to upscale and introduce complexity while preserving the overall coherence and shape of the original cloud.
@@ 7.2 Differentiable Rendering
The propagation of the gradients $\nabla_{\mathcal{R}}\mathcal{L}$ through all volumetric fields requires complex light-path sampling depositing the radiative quantities at every path interactions.

### v3

@@ Abstract
Figure 1: Given a single view ($y$) of a volume ($V$), we reconstruct a volume ($\hat{V}$) from its latent representation ($\theta$) that matches $y$ under the same lighting conditions, resulting in a synthesized view ($\hat{y}$). A differentiable volume renderer ($\mathcal{R}$) is used to optimize physical scene parameters ($\phi$) while simultaneously performing posterior sampling $p(\theta|y;\phi)$, conditioned on the observation, in the latent space of a trained diffusion model $p(\theta)$. Ambiguities due to the absence of information about unseen parts of the volume are reduced by gradually steering the reverse diffusion process toward the most plausible reconstruction under the given view (right section).
@@ 1 Introduction
Differentiable rendering (DR) leverages a rendering process with gradients, making it suitable for recovering shape and optical material parameters from images [25, 70, 37, 11]. DR enables backpropagation of gradients of a loss in image space to the scene parameters, including position, texture, lighting, shape, and other attributes. The challenge increases significantly when these parameters describe complex distributions of volumetric materials, such as clouds, smoke, or fire. In such scenarios, the problem becomes so ill-posed that it requires dozens, if not hundreds, of different views to adequately constrain the object parameters  [44, 45, 27]. It is now widely accepted that reconstructing the internal density distribution of highly dense volumes is nearly impossible due to the high uncertainty in the light scattering process and the presence of vanishing gradients. This limitation can only be alleviated by incorporating prior information during reconstruction.
@@
Our proposed approach addresses these challenges by employing a diffusion prior to guide a Physically-based Differentiable Volume Renderer (PDVR) toward reconstructing a plausible volumetric field. In contrast to previous approaches, our approach includes controlled variations in the diffusion step by considering the gradient of the image loss with respect to the learned latent space representation. This approach steers the reconstruction toward a realistic 3D density distribution, ensuring that the generated structure aligns well with the observed data and maintains realistic spatial consistency.
@@
  • A novel Parametric Diffusion Posterior Sampling (PDPS) technique utilizing a shape-centric prior with a physically-based differentiable volume renderer.
@@ 2 Related Work
NeRFs have been generated initially with MLP-based Scene Representation Networks (SRNs) [55], which have later been used to compactly encode volumetric scalar fields using the emission-absorption optical model  [33, 74]. Alternative to the use of SRNs, adversarial approaches have recently emerged. They use 2D images to stochastically condition the 3D reconstruction using an adversarial loss [53, 4, 16, 43, 87]. In this context, sparse tri-plane volumetric models have been proposed to reduce the memory consumption at improved training efficiency of NeRFs [14, 5]. While NeRF-based approaches usually assume that images of the scene from many different viewpoints exist, recent advancements have shown their potential to also perform single-view reconstruction [4, 69, 53, 39, 76, 29].
@@
Zhang et al.  [80] present a general framework, including volumetric media, for calculating radiance derivatives with respect to changes of scene parameters. This framework has later been extended to make it applicable to path tracing including random sampling [81]. Properties of the differentiation of integrators are analyzed by Zeltner and Monte [77]. Forward mode automatic differentiation [44] for differentiable rendering is nowadays replaced by radiative backpropagation [45] to decrease the required memory, yet at the expense of multiple branches along light paths and quadratic time complexity thereof. Performance increases are achieved by reusing radiances along light paths [67], and by avoiding recursive radiance estimates at scattering locations with dedicated sampling methods for estimating derivatives of volumetric scattering  [46]. For multi-view reconstruction of volumetric fields in the presence of global light transport, singular path sampling in combination with in-scattering relaxation and an exponential moving average shows improved reconstruction fidelity [27]. Under the assumption of an emission-absorption optical model, the “inversion trick” enables fast automatic differentiation for volume reconstruction and transfer function learning  [73]. Physical constraints are combined with self-supervision for the reconstruction of single-scattering flow fields from single-view videos [12].
@@ 4 Method
Figure 2: Top images: Cloudy Dataset – Photorealistic renderings of randomly selected clouds from our dataset, illustrating natural variations and details. Bottom images: Diffusion-based cloud synthesis – Clouds generated with our diffusion model, demonstrating a convincing appearance under realistic lighting conditions and physical parameters.
@@ 4.1 Cloudy - a 3D Clouds Dataset
To add natural randomness and represent diverse distributions of warm columns to the clouds, we apply Perlin noise functions and varied particle emission shapes. Figure 2 (top) shows a random selection of clouds from our dataset, which are rendered under different lighting conditions. The density fields are numerically simulated on regular 3D grids at a resolution of approximately $512\times 256\times 512$.
@@ 4.2 Volume Latent Encoding
We introduce an implicit neural representation for a volume $\mathcal{V}$ defined on the cube $[-1,1]^{3}$, based on a single projection, which we refer to as monoplanar. Unlike previous approaches that use positional feature embeddings like triplane or tensor decomposition, our method involves sampling a window across a single projected axis, centered at the coordinate of interest, to extract the final features.
Let $g:\mathbb{R}^{2}\rightarrow\mathbb{R}^{N}$ be a continuous two-dimensional field of features based on a grid, i.e., $g(x,y)$ returns a $1$-dimensional vector with $N$ sampled values using bicubic interpolation. The vector is structured as another grid $\mathcal{F}$ with domain $[-1,1]$. The function $f(z;\mathcal{F})$ samples $\mathcal{F}$ at positions $z-1+k*\Delta,k\in\{0\dots N-1\},\Delta=2/(N-1)$ using linear interpolation. Sampled positions are constrained to $[-1,1]$. The feature vector $\mathcal{F^{\prime}}$ storing the $N$ interpolated samples is fed into an MLP to produce the final density value, see Figure 3.
@@
In practice, we parameterize $g$ with a coarse grid $\theta$ of size $128\times 128$ and $32$ features. A convolutional upsampler is applied to increase the resolution to $256\times 256\times 64$. Once upsampled, the feature vector at a specific position $(x_{0},y_{0},z_{0})$ is obtained using $g$ and $f$ described earlier.
@@ 4.3 Volume Latent Space
Since cloud formations are equivariant to arbitrary rotations and minor scaling along the $xy$-plane, we apply 14 such operations to the clouds and augment the dataset by these instances. The analog transformations are applied to the latent codes as an initial solution, which is then subsequently refined via optimization. While the transformed latent already represents a plausible volume, the refinement prevents the diffuser from learning patterns that emerge purely from resampling, i.e., due to boundaries and clamping (see the supplementary material for an example). Including the $8$ equivariant transformations (flips and transposes), we obtain a total of $1,000\times 14\times 8$ volume instances for training. Figure 2 (bottom) demonstrates the effectiveness of our diffuser in generating new, unconditional volumetric instances. The ability to produce clouds with realistic shape and interior is demonstrated in Figure 4.
Figure 4: Diffusion Sampling. First column: A cloud from the Cloudy dataset. Subsequent columns show clouds generated by our diffusion model. First row shows the clouds under neutral lighting conditions, demonstrating realistic cloud-like formations. Bottom row shows cross-sectional slices through the volumes, demonstrating realistic interiors of diffused clouds.
@@ 4.4 Parameterized Posterior Sampling
Let us now assume that a proper posterior sampling method $p(\theta|y;\phi)$ is available, meaning that given an observation $y$ and a forward model $y=\mathcal{A}(\theta;\phi)+\eta$, we can draw samples $\theta$ that satisfy the observation. In our case, $\mathcal{A}(\theta;\phi)$ encapsulates both the decoding of the volume from $\theta$ and the rendering depending on $\phi$, i.e., $\mathcal{A}(\theta;\phi):=\mathcal{R}(\mathcal{D}(\theta),\phi)$.
@@
After determining $\hat{\phi}$, the final latent representation $\theta$ can be sampled from the posterior distribution $\theta\sim p(\theta|y;\hat{\phi})$. In addition to the loss in Eq. 2, we can incorporate a regularization term $\mathcal{L}_{\textsc{reg}}(\phi)$ to enforce additional priors on the physical parameters.
@@ 4.5 Optimization
Depending on the complexity of $\mathcal{A}(\theta;\phi)$ with respect to the parameters, it may be advantageous to reuse the same sample $\theta$ for multiple steps in a pass, during the overall optimization. This strategy reduces the need for repeated sampling – to a small number of passes – while still allowing effective updates to $\phi$ over several iterations. This can be particularly useful when $\mathcal{A}(\theta;\phi)$ involves expensive operations or when the gradient propagation is computationally intensive.
@@
Finally, an optional refinement step can be applied, which enforces data consistency [57] before the latent $\theta$ is reused to improve $\phi$ and diffuse for the next step. This is achieved by directly optimizing the latent without any prior supervision. The rationale is that certain features will be preserved, allowing the latent to converge more quickly without constraints. Additionally, if ambiguity arises, it is advantageous for it to be reflected initially in the parameter that is subsequently “cleaned” by the prior. In practice we applied it a few steps around the middle of the process, to avoid early local minima in the beginning and artifacts due to overfitting at the end. The proposed optimization is outlined in Algorithm 1.
@@
$P$triangleright Number of passes
@@
     else
         $\theta_{s}\leftarrow\hat{\theta}_{s}$
return $\phi_{P},\theta_{P}$
@@ 5 Results
In this section, we demonstrate the effectiveness of our method for different use cases. All modules are implemented in Pytorch [48] and Vulkan SDK. Further details are provided in the supplementary material. The code and the Cloudy dataset are publicly available at https://www.github.com/rendervous/cloudy_project.
@@ 5.1 Diffusion Posterior Sampling
In the first experiment, we shed light on the potential of DPS for single-view volume reconstruction. Through this experiment, we do not optimize for any physical parameters affecting the cloud appearance, but solely assess the strength of the volume diffusion prior when used to constrain the differentiable volume renderer.
@@ 5.2 Monoplanar Representation
To assess the quality that is achieved with the proposed monoplanar latent representation, we perform a series of experiments with the monoplanar, triplanar and dense grid representations. All representations use the same number of parameters for the latent, i.e.: Monoplanar $128\times 128\times 32$, Triplanar $3\times 128\times 128\times 11$, and Grid $32\times 32\times 32\times 16$. An upsampler is used in the cases of monoplanar and triplanar representation.
Table  1 shows the average values for each metric across nine reconstructions using clouds from the Cloudy dataset.
@@
While PSNR, RMSE, and MAE consider the full volume at $256\times 128\times 256$ resolution, SSIM [71] considers the center slice. Our proposed monoplanar representation quantitatively outperforms the other state-of-the-art representations in terms of reconstruction fidelity.
@@ 5.3 Super-Resolution
Figure 7: Cloud Super-Resolution. From a cloud on a $32\times 16\times 32$ grid (center), the diffuser reconstructs a density distribution on a $256\times 128\times 256$ grid (right). This process adds fine details and internal structures, demonstrating the model’s ability to upscale and introduce complexity while preserving the overall coherence and shape of the original cloud (left).
5.4 Cloud Recovery from Transmittance Measures
5.5 Comparative Evaluation
Figure 9: Reconstruction comparison. The four leftmost $2\times 2$ images depict views used for reconstruction and testing. The number on the label indicates if 1 or 3 images were used for the reconstruction. The reconstruction time is reported in minutes (top), along with the LPIPS [82] metric value (bottom), which quantifies the perceptual similarity between the ground truth and the synthesized views.
Metric DRT1 DRT3 SPS1 SPS3 DPS1 DPS3 DPS3x3
[…]
Time 00:30:40 00:40:58 00:31:44 00:33:54 00:03:44 00:04:23 00:14:47
Table 2: Quality comparison of DRT, SPS and DPS (ours) using one and three views for reconstruction. The table shows average values over $32$ test cases, each constructed using clouds, materials, cameras, and environment settings sampled from 16 unseen clouds, 3 distinct cloud materials, 7 different environments, and 5 sets of camera poses.
To compare our novel DPS approach with previous methods for reconstructing 3D clouds from images, we evaluate DPS alongside Differentiable Ratio-Tracking (DRT) [46] and Singular Path Sampling (SPS) [27]. Since both DRT and SPS require multiple views to achieve accurate results, we tested with one and three images for the reconstructions.
We evaluate DPS under three different settings: (1) using only a single view (DPS1), (2) using all three views (DPS3), and (3) performing three restarts of the diffusion from a noisy version of a previously reconstructed latent (DPS3x3). The last setting aligns with diffuse-denoise strategies, progressively adjusting the initial noise toward the observed data to improve guidance stability. Results are shown in Fig. 9 and summarized in Table 2.
5.6 Recovering Light Conditions
Figure 10: Recovering $\phi$. Top: Reconstructions using parameterized DPS under two scenarios – when the background radiance is unknown (Background), and when the entire lighting condition is unknown (Environment). Bottom: Evolution of the recovered background (top) and environment (bottom). Final column shows the lighting condition used to render the test views.
@@ 7.2 Differentiable Rendering
The propagation of the gradients $\nabla_{\mathcal{R}}\mathcal{L}$ through all volumetric fields requires complex light-path sampling depositing the radiative quantities at every path interaction.